The perimeter of a rectangle is 210m. If the length of this rectangle is three-fourths of its width, its dimensions are? I attempted solving this by using the equation P= 2L + 2W; and attempting to solve for the length, etc. I'm still stumped with the 3/4.
Thanks

Answer by checkley79(3341)

L=3/4W
2L+2W=210
2(3W/4)+2W=210
6W/4+2W=210
(6W+4*2M)/4=210
(6W+8W)/4=210
14W/4=210
14W=210*4
14W=840
W=840/14
W=60 ANS.
L=(3/4)60
L=45 ANS.
PROOF:
2*60+2*45=210
120+90=210
210=210